All Systems Operational

RIPE Database Operational
90 days ago
99.92 % uptime
Today
Email to/from the RIPE NCC Operational
90 days ago
99.7 % uptime
Today
K-root/DNS ? Operational
90 days ago
100.0 % uptime
Today
LIR (Member) Portal Operational
90 days ago
99.88 % uptime
Today
RIPE NCC Access Operational
90 days ago
99.88 % uptime
Today
RPKI Operational
90 days ago
99.97 % uptime
Today
RPKI Dashboard Operational
90 days ago
99.88 % uptime
Today
RRDP Repository Operational
90 days ago
100.0 % uptime
Today
rsync Repository Operational
90 days ago
100.0 % uptime
Today
Publication as a Service (API endpoints) ? Operational
90 days ago
100.0 % uptime
Today
www.ripe.net Operational
90 days ago
100.0 % uptime
Today
Reverse DNS Operational
90 days ago
100.0 % uptime
Today
Non-Critical Services Operational
90 days ago
99.95 % uptime
Today
Billing & Payments Operational
RIPEstat Operational
DNSMON Operational
Mailing Lists Operational
RIPE NCC Academy Operational
RIPE NCC Certified Professionals Operational
RIPE Atlas Operational
RIPE Labs Operational
RIS (Routing Information Service) Operational
Meeting Software Operational
Pilot (RPKI) ? Operational
90 days ago
100.0 % uptime
Today
ftp.ripe.net ? Operational
90 days ago
99.91 % uptime
Today
Operational
Degraded Performance
Partial Outage
Major Outage
Maintenance
Major outage
Partial outage
No downtime recorded on this day.
No data exists for this day.
had a major outage.
had a partial outage.
Nov 30, 2025

No incidents reported today.

Nov 29, 2025

No incidents reported.

Nov 28, 2025
Resolved - The link is fully restored and services are stable. A post‑mortem will be published shortly.
Nov 28, 12:57 CET
Update - The link has been fully restored, and the firewall cluster has recovered successfully. All systems are now operating normally, but we will keep monitoring them for the next few hours while we wait to receive a root‑cause explanation from the dark fibre vendor.
Nov 28, 11:23 CET
Update - We are awaiting further information from our dark fibre vendor.
Nov 27, 16:40 CET
Update - We identified the likely underlying cause. A flapping link on one of our dark fibre connections between data centre locations disrupted our firewall cluster as well as other network components. The link subsequently failed completely.

We are currently operating with reduced redundancy. An engineer from our dark fibre vendor is en route to verify the physical connection. We continue to monitor for any further issues.

Nov 27, 14:21 CET
Update - We are continuing to monitor for any further issues.
Nov 27, 13:01 CET
Update - The issue, caused by a network outage, has reoccurred. We are continuing to investigate this issue.
Nov 27, 12:00 CET
Update - The issue, caused by a network outage, has reoccurred. We are continuing to investigate this issue.
Nov 27, 11:57 CET
Monitoring - Connectivity has been restored, and all services are operational. We are currently investigating the root cause.
Nov 27, 11:29 CET
Update - We are continuing to investigate this issue.
Nov 27, 11:21 CET
Investigating - We are currently investigating this issue.
Nov 27, 11:19 CET
Resolved - This incident has been resolved.
Nov 28, 11:02 CET
Monitoring - Probes are now operating normally, and measurement delays have fallen to normal levels.
Nov 27, 15:02 CET
Identified - The issue, caused by a network outage, has reoccurred. Delays are increasing again. We are working on a solution.
Nov 27, 11:52 CET
Monitoring - Probes are now operating normally, and measurement delays have fallen to normal levels.
Nov 27, 11:33 CET
Investigating - We are investigating an issue preventing probes from receiving new measurement instructions and from delivering results.
Nov 27, 11:11 CET
Resolved - This issue was related to recovering all connections after the datacenter interconnect had partly failed
https://status.ripe.net/incidents/jmm1527lcj67

Nov 28, 09:58 CET
Monitoring - Connectivity has been restored, and all services are operational. We are currently investigating the root cause.
Nov 28, 09:49 CET
Investigating - We are currently investigating this issue.
Nov 28, 09:43 CET
Nov 27, 2025
Nov 26, 2025
Completed - The scheduled maintenance has been completed.
Nov 26, 16:00 CET
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Nov 26, 14:00 CET
Scheduled - We will be upgrading the RIPE Atlas controllers, which will cause all probes and anchors to disconnect and remain disconnected for a short period, expected to be around 30 minutes. During this time probes will continue to carry out measurements, but there will be delays in how quickly the results will be available.
Nov 25, 16:44 CET
Nov 25, 2025
Resolved - BGP sessions are re-established and data collection has resumed as of 24 November 20:45 UTC . Data processing has caught up. This incident has been resolved.
Nov 25, 08:25 CET
Monitoring - The connectivity of RRC15 has been restored.
BGP sessions are established and the MRT files are being produced.
We are monitoring the situation.

Nov 25, 00:12 CET
Identified - Since 23 November, 07:00 UTC, we are not receiving BGP data on RIS route collector rrc15 due to a connectivity issue between the collector and the internet exchange. This means the BGP sessions are down. As a result, the MRT update and bview files are close to empty.

We are working on restoring connectivity. Data collection will resume when connectivity is restored.

Nov 24, 11:12 CET
Completed - The scheduled maintenance has been completed.
Nov 25, 08:15 CET
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Nov 25, 08:00 CET
Scheduled - We plan to perform a switchover of the RIPE database primary instance, so we can perform patching on the server. We don't expect any downtime for Whois queries or updates.
Nov 17, 11:14 CET
Nov 24, 2025
Nov 23, 2025

No incidents reported.

Nov 22, 2025

No incidents reported.

Nov 21, 2025
Resolved - Previous results have been processed, the data flow is back to normal.
Nov 21, 18:05 CET
Monitoring - We have restored the pipelines, older results are being backfilled into BigQuery now.
Nov 20, 16:40 CET
Identified - We are working on a solution to restore the data pipelines
Nov 20, 16:00 CET
Investigating - The pipeline to ingest RIPE Atlas results into Google BigQuery is down. We're investigating the cause for this event.
Nov 20, 08:51 CET
Nov 20, 2025
Nov 19, 2025
Completed - The scheduled maintenance has been completed.
Nov 19, 21:43 CET
Update - Scheduled maintenance is still in progress. We will provide updates as necessary.
Nov 19, 21:15 CET
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Nov 19, 21:15 CET
Scheduled - RPKI Dashboard is unavailable due to planned maintenance.
Nov 11, 20:19 CET
Resolved - This incident has been resolved. A configuration change allowed data processing to catch up.
Nov 19, 07:24 CET
Update - We are continuing to work on this issue. At the moment, data is available through 2025-11-14.
Nov 18, 09:17 CET
Identified - Our backend system indexing BGP updates has not updated since 2025-11-11 20:32 UTC. An earlier disruption caused a backlog of data, which changed the distribution of the data we need to process. This in turn caused problems in our processing pipeline.

As a result the data in the RIPEstat bgp-state, bgp-updates, and bgplay endpoints has not updated since this time.

We are working to resolve the processing issues and clear the backlog. Due to the low impact of this issue, resolution may continue after the weekend.

Nov 14, 16:42 CET
Nov 18, 2025
Completed - The scheduled maintenance has been completed.
Nov 18, 12:09 CET
Verifying - Verification is currently underway for the maintenance items.
Nov 18, 11:43 CET
In progress - Scheduled maintenance is currently in progress. We will provide updates as necessary.
Nov 18, 10:09 CET
Scheduled - We will carry out planned maintenance on the Atlas Kafka cluster. During this time, measurements will continue running but new results may not be available.
Nov 18, 10:08 CET
Nov 17, 2025

No incidents reported.

Nov 16, 2025

No incidents reported.